# MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark

### v1

@@ A.4.1 Verilog/SystemVerilog
A.4.2 High-Level Synthesis
@@ A.4.4 Chisel
(b) HLS: Structural Complexity
(c) SystemVerilog: Line Number
@@
Figure 10: EM performance for HLS and SystemVerilog. Top-to-bottom: line number then structural complexity.
(a) Chisel: Line Number
(b) Chisel: Structural Complexity
(c) VHDL: Line Number
(d) VHDL: Structural Complexity
Figure 11: EM performance for Chisel and VHDL. Top-to-bottom: line number then structural complexity.
A.5 GraphCoder Modifications
The original implementation of GraphCoder is designed for left-to-right completions using only the left context. To fit our objective of fill-in-the-middle completions, we modified GraphCoder such that the right context is also preserved and included during prompt assembly. As the graph builder and the retriever operate independently of the left and right contexts, this has no other effect on the functionality of GraphCoder.
Furthermore, the original implementation of GraphCoder included helper functions in utils/ccg.py for traversing the CST that only supported Java and Python. To extend support to hardware languages, we added helper functions to handle each language’s unique syntax and identify control-flow and data-dependence. Other minor changes were made in various locations to include HDL languages during tree-sitter Parser language selection and code file filtering.
@@ A.5 GraphCoder Modifications
Table 8: Retriever hit@k on HDL languages and Python
A.6 CodeBLEU
@@
We adopt a practical, language-aware variant of CodeBLEU to evaluate hardware code generation across Chisel, SystemVerilog, VHDL, and HLS (C/C++). The final score is computed as a weighted combination of four complementary components:
@@ Discussion.
A.7 Compilation rate calculation.
@@
A.8 Detailed Performance on line numbers and structural levels
We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT-5) varies with the number of lines to be completed and structural complexity for different hardware languages in Figures 10 and 11.
A.9 Detailed Performance on hardware semantic levels.
@@ A.8 Detailed Performance on line numbers and structural levels
Figure 12: Per-category performance on Chisel. Metrics are reported in terms of EM, ES, and BLEU.
@@
Figure 14: Per-category performance on HLS code completion.
@@
Figure 15: Per-category performance on VHDL across different evaluation metrics.
@@
A.10 Code Completion Examples
For each hardware description language, we select two representative repositories as examples to illustrate cross-file context, in-file context, the selected completion target, and the corresponding ground truth in the following figures.
@@ A.10 Code Completion Examples
Verilog AXI-SDCard-High-Speed-Controller input logic[2:0] axilite_arprot, .waddr(waddr), .wdata(wdata), .we(we)... (Hallucinated Instantiation)
@@
Verilog SystemVerilogAssertions logic gnt3; // Write assertions for a simple Round Robin... (Context Reset)
@@
HLS hls-spmv L1: for (int i = 0; ... void spmv(int rowPtr[NUM_ROWS+1]... (Function Definition)
HLS libllsm FP_TYPE* x = calloc(ns... FP_TYPE* llsm_envelope(FP_TYPE* spectrum... (Function Definition)
@@
VHDL cpu_for_nscscc2022_single C_HAS_WE => 1, C_MEM_... --- FILE: dcache_tag.vhd (END) --- (Premature EOF)
VHDL FPGA-I2C-Minion if ENDSIM = false then... i2c_wait_quarter_clock; scl_pre_spike <= ’Z’; (Unrelated Logic)
Chisel rjrouter class Lfsr16Test extends... <|fim_middle|> (Model Failure Token)
Table 9: Comparison of Ground Truth vs. Pre-Training Failures across Hardware Languages.

### v2

@@ A.5.1 Verilog/SystemVerilog
A.5.2 High-Level Synthesis
@@ A.5.4 Chisel
(b) SystemVerilog: Line Number
(c) HLS: Structural Complexity
@@
(e) Chisel: Line Number
(f) VHDL: Line Number
(g) Chisel: Structural Complexity
(h) VHDL: Structural Complexity
Figure 10: EM performance across languages. Rows correspond to line number (rows 1 and 3) and structural complexity (rows 2 and 4).
A.6 GraphCoder Modifications and Limitations
The original implementation of GraphCoder is designed for left-to-right completions using only the left context. To evaluate its performance on fill-in-the-middle tasks, we modified GraphCoder such that the right context is also preserved and included during prompt assembly. As the graph builder and the retriever operate independently of the left and right contexts, this has no other effect on the functionality of GraphCoder.
Furthermore, the original implementation of GraphCoder included helper functions in utils/ccg.py for traversing the CST that only supported Java and Python. To extend support to hardware languages, we added helper functions to handle each language’s unique syntax and identify control-flow and data dependence. Other minor changes were made in various locations to include HDL languages during tree-sitter parser language selection and code file filtering.
@@ A.6 GraphCoder Modifications and Limitations
Further analysis of retrieval results have revealed properties of HDL languages that rendered subtle design features in GraphCoder to become liabilities to effective retrieval. Using HLS as example, we analyze the retrieval results given by GraphCoder and RepoCoder and discover two key differences in the characteristics of retrieved code snippets.
Absence of Function Declarations: We found that 16.9% of retrieved code snippets from RepoCoder contains HLS function definitions, while the number drops to 1.8% for GraphCoder. This arises from a design in the official implementation of GraphCoder: In Line 79 of utils/slicing.py, GraphCoder explicitly filters out nodes with definition in their type names, likely to focus on executable statements. However, it directly causes many function signatures to be excluded and affects HLS retrieval because key synthesis constraints (types, interfaces, and pragmas that shape parallelism) are often encoded in definitions and headers rather than in nearby data/control-flow statements.
Absence of HLS Constructs: As there exists no implementation of HLS grammar for tree-sitter, tree-sitter-cpp was used to extend support to HLS. However, the C++ grammar of tree-sitter does not process HLS-specific constructs, causing them to be left out of retrieval. For instance, 4.2% of snippets given by RepoCoder contains HLS-specific directives, while the rate for GraphCoder is only 0.7%. This causes many information-carrying synthesis constraints to be systematically omitted from GraphCoder’s retrieved context. In HLS, pragmas and vendor-specific types are not merely stylistic annotations: they encode parallelization, memory partitioning, and interface protocols that directly affect correctness and implementability.
Furthermore, we notice a weakness in code snippets captured by GraphCoder that is present across all HDL languages included in MHRC-Bench: code snippet length divergence. We calculate the mean and standard deviation in character length for code snippets across RepoCoder and GraphCoder, the results of which are reported in Table 9. The figures indicate that GraphCoder’s retrieval snippets have half the mean but triple the standard deviation of the snippets produced by RepoCoder and suggests an instability in GraphCoder’s processing of CST structures of HDL languages.
Table 9: Snippet length statistics (characters) over the center 90% of retrieved snippets.
Retriever # Snippets (center 90%) Mean Length (chars) Std. Dev. (chars)
GraphCoder 1595 2868.10 3223.60
RepoCoder 1596 4536.98 1060.49
Table 10: Retriever hit@k on HDL languages and Python
@@
A.7 CodeBLEU
@@
We adopt a practical, language-aware variant of CodeBLEU to evaluate hardware code generation across Chisel, SystemVerilog, VHDL, and HLS (C/C++). The final score is computed as a weighted combination of four complementary components. These weights and CodeBLEU calculation are set according to Ren et al. (2020).
@@ Discussion.
A.8 Compilation rate calculation.
@@
A.9 Detailed Performance on line numbers and structural levels
We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT-5) varies with the number of lines to be completed and structural complexity for different hardware languages in Figure 10.
A.10 Detailed Performance on hardware semantic levels.
@@ A.9 Detailed Performance on line numbers and structural levels
Figure 11: Per-category performance on Chisel across EM, ES, and BLEU.
@@
Figure 13: Per-category performance on HLS across EM, ES, and BLEU..
@@
Figure 14: Per-category performance on VHDL across EM, ES, and BLEU.
@@
A.11 Code Completion Examples
For each hardware description language, we select two representative repositories as examples to illustrate cross-file context, in-file context, the selected completion target, and the corresponding ground truth in Figures 15 to 22.
Table 11: Comparison of Ground Truth vs. Pre-Training Failures across Hardware Languages
(Examples extracted from Gemini 3 Pro output).
@@ A.11 Code Completion Examples
Verilog AXI-SDCard-High-Speed-Controller input logic[2:0] axilite_arprot, .waddr(waddr), .wdata(wdata), .we(we)... (Instantiation Hallucination)
@@
Verilog SystemVerilogAssertions logic gnt3; // Write assertions for a simple Round Robin... (Non-code Content Hallucination)
@@
HLS hls-spmv L1: for (int i = 0; ... void spmv(int rowPtr[NUM_ROWS+1]... (Function Definition Hallucination)
HLS libllsm FP_TYPE* x = calloc(ns... FP_TYPE* llsm_envelope(FP_TYPE* spectrum... (Function Definition Hallucination)
@@
VHDL cpu_for_nscscc2022_single C_HAS_WE => 1, C_MEM_... --- FILE: dcache_tag.vhd (END) --- (Non-code Content Hallucination)
VHDL FPGA-I2C-Minion if ENDSIM = false then... i2c_wait_quarter_clock; scl_pre_spike <= ’Z’; (Logic Hallucination)
Chisel rjrouter class Lfsr16Test extends... <|fim_middle|> (Non-code Content Hallucination)